# Are Existing Road Design Guidelines Suitable for Autonomous Vehicles?

### v1

@@ 0 Apollo - 7 9 5 7 7
Methods TrashFuzz (Ours) MSF [11] ADSDoS [45] DeepBillboard [23] Dirty Road [14] Car Patch [15] Frustum [12] AttackZone [16]
Patch? $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$
Shape? $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Placement? $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
Special? $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$
White-B? $\times$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$
Grey-B? $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$
Black-B? $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\times$
RQ5: How difficult is it to carry out $\mathtt{TrashFuzz}$ attacks? To answer this question, we propose a comparison between the prerequisites of existing attacks and those of $\mathtt{TrashFuzz}$. The results are shown in Table VIII. Within the table, the categories of ‘Patch?’, ‘Shape?’, ‘Placement?’, and ‘Special?’ respectively assess whether the listed attack necessitates a specialised adversarial patch, alters the shape of the adversarial object, designates specific locations for placing adversarial elements, or requires unique equipment unlikely to be found on the roadside (e.g., commercial drones, digital LiDAR spoofers, and projectors). Furthermore, ‘White-B?’, ‘Grey-B?’ and ‘Black-B?’ categorise the type of attack. We mark $\surd$ in the table to denote alignment with each criterion. For example, the MSF-Attack [11] alters the shape of objects (e.g. traffic cones) to make them invisible to the camera and LiDAR detection algorithm. Hence, Shape? and Placement are marked $\surd$.
As can be seen from Table VIII, our attack necessitates neither adversarial patches nor intricate shape modifications as in the previous works. Hence, we do not have concerns regarding the printability of adversarial objects. Moreover, our attack operates as a black box attack, making it applicable to a broader range of situations. While LiDAR Seg [46] and ADS-DoS [45] share similar prerequisites with $\mathtt{TrashFuzz}$, LiDAR Seg focuses on deceiving the LiDAR detection algorithm without considering the camera detection algorithm and ADS-DoS is a grey-box attack which requires knowledge of the planning module of the ADS. In contrast, $\mathtt{TrashFuzz}$ takes the entire perception system into consideration and is a black box attack. In summary, when compared to existing attacks, it can be concluded that $\mathtt{TrashFuzz}$ is relatively easy to conduct.
Threats to Validity. Given the nature of simulation-based testing, it is important to acknowledge the presence of potential validity threats to the identified issues. For instance, the ADS consumes a substantial amount of computational resources, and limitations in computational resources can introduce latency into the ADS. To address this challenge, we employ a solution by running the ADS and simulator on two separate machines connected by an Ethernet cable. During the experiment, we ensure memory cleanliness and exclude any concurrent tasks. As a result, we consistently achieve an average latency of less than 0.1 seconds for all the Apollo modules during simulations. Furthermore, we replicate the identification of issues at least three times to ensure that these issues are not the result of random latency. Despite taking these measures, in general, we cannot definitively rule out the possibility that a discovered issue might be influenced by the running environment. Nevertheless, identifying such problems remains valuable for enhancing the overall system.
5 Related Work
Attacks for AV. ADSs are complex and security-critical systems, but ensuring their resilience against atacks is difficult. Two categories of works have introduced several attacks to analyse security issues of the ADS, which we review below.
The first category of research focuses on attacking the ADS based on adversarial patches. In several attacks [23, 24, 25], patches have been strategically placed on roadside billboards to influence the behaviour of the ADS. In another, a simple flat dirty patch on the road surface can disrupt the Automated Lane Centering of an ADS, causing it to deviate from the intended steering angle [14]. Similarly, the presence of an adversarial patch on a vehicle in front can interfere with the depth estimation algorithm of the ADS, potentially leading to collisions [15]. Finally, the use of projectors to create ‘shadows’ can introduce errors into the segmentation algorithm of the ADS, causing deviations in the bounding box of detected objects from the ground truth [16].
The second category of attacks explores how to attack ADSs based on the placement of objects. For example, MAF-ADV [11] manipulates the shape of an object, such as a traffic cone, making it invisible to both cameras and LiDAR object detection algorithms. Then, the attacker can place the adversarial object in front of the ADS to cause a collision. In the research by Zhu et al. [13, 46], objects such as cardboard, road signs, or commercial drones are strategically positioned at specific locations to reflect laser beams, generating misleading points that disrupt the LiDAR object detection and segmentation algorithms. The frustum attack [12] exploits the camera’s blind spots, utilising these regions to influence the LiDAR segmentation algorithm and consequently affecting the fusion of camera and LiDAR results. Wan et al. [45] strategically place objects such as traffic cones to provoke overly conservative behaviour within the planning module of the ADS. Furthermore, unconventional ‘objects’ like invisible lights are employed to disrupt the camera perception [47].
The aforementioned attacks rely on conspicuous patches or objects with non-standard shapes. Additionally, these adversarial attacks are mainly white-box attacks, which may be unrealistic in practice. With $\mathtt{TrashFuzz}$, we present a black box attack that has the capability to induce hesitation or trigger aggressive driving behaviour in a state-of-the-art ADS Apollo. Furthermore, these behaviours can be induced by placements of objects that satisfy guidelines for road design.
Critical Scenario Generation. A number of works explore the safety and efficiency issues of ADS by generating critical scenarios. Part of these scenario-generation approaches centre around altering the behaviour of background vehicles to cause collisions. For instance, AVFuzzer [48] and DoppelTest [49] employ a genetic algorithm-based fuzzing engine to optimise scenarios, primarily focusing on minimizing the distance from other background vehicles. Similarly, AutoFuzz [50] proposes a constrained neural network evolutionary search approach to guide AVs into scenarios resulting in collisions or veering off the road. NADE [34] initially generates scenarios mimicking the natural driving behaviour of background vehicles but introduces deviations in their driving behaviour at key points to create adversarial scenarios likely to lead to collisions. CRISCO [51] assigns participants to execute impactful behaviors aimed at challenging the ADS, ultimately leading to collisions involving AVs. AVUnit [52] introduces a scenario description language that characterises the driving behaviour of background vehicles, pedestrians, and weather conditions, proposing a fuzzing algorithm based on encoding scenarios described in this language. DeepCollision [53] employs a Reinforcement Learning algorithm to manipulate the driving environment, directing AVs toward collision scenarios. MOSAT [54] uses a multi-objective genetic algorithm to search for scenarios that cause collisions. Similarly, Rule-based Searching [55], Evolutionary-Algorithm-based Generation [56], and CMTS [57] all manipulate the driving behaviours of background vehicles to induce collision scenarios.
In addition to collision avoidance, more elaborate optimisation objectives have been explored as well. LawBreaker [17] uses traffic laws as specifications and presents a coverage-based fuzzing algorithm to trigger various traffic law violations by modifying the driving behaviours of background vehicles. ABLE [58] employs active learning to efficiently extend the coverage of traffic laws. TARGET [59] uses GPT-3 to analyse traffic laws, thereby enabling the formulation of specifications for generating test scenarios. Althoff et al. [60] aim to minimise the passable area by modifying the driving behaviours of background vehicles. Metamorphic Testing of Driverless Cars [61] focuses on checking the relationships between the inputs and outputs across multiple executions of the ADS. The Baidu Group introduces a coverage-based feedback mechanism in their work [62], where the criterion is the extent of coverage of the driving area on the map. In other words, achieving a broader coverage of the driving area signifies a more favourable scenario. Hauer et al. [63] investigate the automated and manual modification of the driving behaviours of the background vehicles to cover all categories, framing this task as a Coupon Collector’s problem. MORLOT [64] proposes a method of many-Objective Reinforcement Learning for Online Testing of ADS, guiding AVs to violate 11 safety metrics, e.g. red light and collision infractions.
These works, however, primarily concentrate on mutating the driving behaviour of background vehicles, often overlooking the impact of object placements on ADS behaviour due to their unique perception systems. In this work, we investigate the influence of roadside objects on ADS, revealing that these objects can also significantly affect ADS driving behaviour, even when their placement follows the road design guidelines that characterise our natural scenarios.
6 Conclusion
In this paper, we introduced $\mathtt{TrashFuzz}$, the first black box ADS testing approach based on generating natural (patch-free) adversarial scenarios that respect regulatory and company guidelines for the placement of roadside objects. Our approach utilises a greedy gradient-based fuzzing algorithm that places these objects so as to exploit vulnerabilities in the ADS’s perception system and induce traffic law violations. We implemented $\mathtt{TrashFuzz}$ for the Apollo ADS and LGSVL simulator, and assembled an extensive library of everyday roadside objects using assets from the Unity Asset Store. We found that our approach could uncover natural scenarios that induced Apollo to violate 15 different Chinese traffic laws, including one in which an otherwise benign placement of bins overloaded Apollo’s perception system and caused it to misperceive an adjacent traffic light as green. These results highlight a current lack of resilience in a key ADS, and perhaps suggest guidelines for the placement of roadside objects may need to take into account the presence of AVs on the road. For example, regulating the shape and size of roadside trash bins may help prevent misclassifications, ensuring they are not mistaken for pedestrians or vehicles from any angle.

### v3

@@ V-B Practicality of TrashFuzz Attacks
Method Ours MSF[5] ADSDoS[66] DeepBB[63] DirtyRd[47] CarPatch[10] Frustum[17] AttackZ[35] Patch? $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$ Shape? $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ Placement? $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ Special? $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$ White-box? $\times$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$ Grey-box? $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ Black-box? $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\times$
As shown in Table VIII, TrashFuzz does not require adversarial patches or shape modifications to objects, avoiding the practical challenges associated with designing and deploying visually conspicuous artefacts. Instead, the attack operates by manipulating the placement of ordinary roadside objects while respecting existing road-design guidelines. Furthermore, TrashFuzz operates entirely in a black-box setting, without requiring access to the internal models of the ADS. Some prior attacks share certain characteristics (e.g., placement-based manipulation), but may require additional assumptions such as knowledge of specific system components or specialised equipment. Overall, TrashFuzz relies only on common roadside objects and black-box access to the ADS, suggesting that such scenarios are plausible to realise in practice.
V-C Limitations and Future Work
Recent AV stacks increasingly incorporate multimodal foundation models (e.g., GPT-4o [38, 39]) that jointly process vision, language, and sensor signals [60, 25]. Extending TrashFuzz to such systems requires moving beyond physical object placement to semantically grounded perturbations, e.g., misleading textual cues or scene-level affordance manipulations that bias multimodal decision-making. Language-guided scene optimisation is a promising direction for testing next-generation AVs.
Our evaluation focuses on object-placement-based scenario generation and does not include reinforcement-learning approaches (e.g., CRASH [27]), which offer a complementary direction. While our work centres on vulnerability discovery, natural extensions include exploring defences such as perception filtering or refining road-design rules (e.g., VisionGuard [18]). We also focus on static roadside objects; extending the framework to dynamic elements such as mobile billboards [43] would further broaden the evaluation.
VI Related Work
Attacks for AVs. ADSs are complex and security-critical systems, yet difficult to harden against adversarial behaviours. Existing work largely falls into two categories.
Adversarial-patch attacks. Several studies place adversarial patches on roadside billboards to mislead perception [63, 42, 43], or apply inconspicuous patches on the road surface to disrupt lane-centering [47]. Patches placed on leading vehicles can corrupt depth estimation [10], while projector-based “shadow” patterns induce segmentation errors [35]. These attacks require suspicious-looking artifacts and are typically white-box, limiting practical deployability.
Object-placement attacks. Approaches in this category manipulate the placement or shapes of roadside objects. MAF-ADV makes traffic cones invisible to camera and LiDAR, then positions them to trigger collisions [5]. Other attacks position reflective objects (e.g., cardboard, signs, drones) to inject false LiDAR points [67, 66]. The frustum attack exploits camera blind spots to mislead multi-sensor fusion [17]. Wan et al. [57] and Wang et al. [58] show that cones, lights, and other objects can induce overly conservative or incorrect behaviours.
These attacks generally rely on conspicuous or non-standard objects and assume white-box access. In contrast, TrashFuzz performs a black-box attack using only everyday objects placed in guideline-compliant (“natural”) locations, yet still induces hesitation or aggressive driving in Apollo.
Critical Scenario Generation. Several approaches generate adversarial or safety-critical scenarios by perturbing background-vehicle behaviours. Genetic-algorithm fuzzers (e.g., AV-Fuzzer [31], DoppelTest [24]) and evolutionary searches such as AutoFuzz [62] optimise scenarios for collisions or off-road outcomes. NADE [15] perturbs naturalistic trajectories, while CRISCO [53] and MOSAT [52] assign impactful behaviours or solve multi-objective formulations for inducing failures. Reinforcement-learning approaches (e.g., DeepCollision [32]) similarly learn to steer scenes toward collisions.
Recent work has also explored objectives beyond inducing collisions [64]. LawBreaker [50] and ABLE [61] target traffic-law coverage; TARGET [11] uses LLMs to extract formalised specifications; other methods evaluate metamorphic consistency [65], derive metamorphic relations from traffic rules to generate transformed road images [12], maximise map-area coverage [23], or frame behavioural diversity as coupon-collector problems [20]. MORLOT [19] guides AVs to violate multiple safety metrics. Recent behaviour-centric methods such as Tian et al. [53] and BehAVExplor [9] extract influential interaction patterns or maximise behavioural diversity to elicit failures—approaches that complement our object-centric, regulation-constrained focus.
Overall, these scenario-generation methods generally mutate agent behaviours, overlooking the effects of naturally-placed roadside objects focused on by TrashFuzz.
VII Conclusion
We introduced TrashFuzz, a black-box approach for testing ADS perception by generating natural, patch-free adversarial scenarios that comply with regulatory and company guidelines for roadside object placement. Using a greedy gradient-based fuzzing algorithm, TrashFuzz identifies placements of everyday objects that exploit perception vulnerabilities and trigger traffic-law violations.
We implemented TrashFuzz for the Apollo ADS with the LGSVL simulator and built a diverse library of realistic roadside objects. Our evaluation uncovered natural scenarios causing Apollo to violate 15 Chinese traffic laws, including a case where a benign bin layout overloaded perception and triggered a traffic-light misclassification.
These results expose robustness gaps in modern ADS perception and suggest that existing roadside-object guidelines may need revision—e.g., standardising the appearance or dimensions of common objects such as trash bins—to better support AV perception systems.
Acknowledgement
This work is partially supported by DTC project DTC-IGC-08.